# On the electronic normal modes of the Meyer–Miller–Stock–Thoss representation of non-adiabatic dynamics

### v1

@@ Abstract
Accurate and efficient simulation of non-adiabatic dynamics is highly desirable for understanding charge and energy transfer in complex systems. A key criterion for obtaining an accurate method is conservation of the Quantum Boltzmann Distribution (QBD). For a single surface, Matsubara dynamics is known to conserve the QBD, as a consequence of truncating the dynamics in the higher normal modes of the imaginary-time path integral. Recently, a Non-Adiabatic Matsubara (NA-Mats) dynamics has been proposed (J. Chem. Phys., 2021, 154, 124124) which truncates in the normal modes of the nuclear variables but not in the electronic variables, which are described with the Meyer–Miller–Stock–Thoss (MMST) representation. Surprisingly, this NA-Mats method does not appear to conserve the QBD for a general system. This poses the question of the effect of truncating the higher normal modes of the electronic variables in the MMST representation. In this article, we present what we believe is the first study of electronic normal modes of the MMST representation. We find that observables are not usually a function of a finite number of normal modes and that the higher normal modes are not constrained by the distribution, unlike in conventional nuclear normal modes. While for an ensemble of trajectories, truncation in MMST normal modes appears to numerically conserve the QBD, this is not the case for a single trajectory. When considering a correlation function, all MMST normal modes are required to obtain reasonable dynamics. Overall, this suggests that MMST variables are not optimal for obtaining an accurate, QBD conserving nonadiabatic dynamics method.
@@ I Introduction
Many of the interesting and complex phenomena observed from charge and energy-transfer mechanisms occur via non-adiabatic dynamics, where the electronic and nuclear degrees of freedom are coupled. Examples include photoisomerism, ultrafast relaxation via intersystem crossing in DNA or charge carriers in solar cell materials, and electron transfers in various photosynthetic systems and catalytic materials.Polli et al. (2010); Wan et al. (1999); Hammes-Schiffer and Stuchebrukhov (2010); Marcus and Sutin (1985); Zhu and Podzorov (2015); Cheng and Fleming (2009); Domcke and Yarkony (2012) As these processes often occur on very short timescales, it can be experimentally challenging to understand the dynamics.
Hence, many theoretical methods have been developed to simulate non-adiabatic dynamics. These range from full quantum methods such as Multi-Configurational Time-Dependent Hartree-Fock (MCTDH) and wavepacket propagation,Beck et al. (2000); Van Haeften, Ash, and Worth (2023); Wang and Thoss (2003); Shin and Metiu (1996) to more approximate semiclassical methods including initial-value representations (IVR),Stock and Thoss (1997); Miller (1970); Wang, Sun, and Miller (1998); Sun, Wang, and Miller (1998a, b); Church et al. (2018); Ananth, Venkataraman, and Miller (2007) non-adiabatic ring-polymer molecular dynamics (RPMD) and their derivatives.Richardson and Thoss (2013); Ananth (2013); Hele (2011); Hele et al. (2015a); Cao and Voth (1993, 1994a, 1994b, 1994c); Rossi, Ceriotti, and Manolopoulos (2014) Another useful class of methods are mixed quantum-classical methods, including surface hopping, Ehrenfest dynamics, and mapping-based methods.Kapral (2016); Tully and Preston (1971); Tully (2012); Shakib and Huo (2017); Shalashilin (2011); Zimmermann and Vaníček (2014); Meyer and Miller (1979); Stock and Thoss (1997, 2005); Runeson and Richardson (2019); Mannouch and Richardson (2023) These methods often have a good balance between computational cost and accuracy.
@@
When considering the development of a non-adiabatic dynamics method, it would be highly desirable if the method satisfied the following criteria:
@@
While many of the existing methods satisfy several of these criteria, as far as we are aware, none satisfy all of them simultaneously and exactly for a general quantum system.
The flow for method development has typically been creating a method and then testing for the properties. To ensure that the method satisfies the desired criteria, we suggest starting with understanding how to obtain these properties and then including this in the method development.
@@
Conservation of the QBD is a particularly interesting criterion to investigate, as for single-surface dynamics, Matsubara dynamics conserves the QBD.Willatt (2017); Hele et al. (2015a) Matsubara dynamics achieves this by starting with an (exact quantum) Generalized Kubo Transform,Hele and Althorpe (2013a, b); Althorpe and Hele (2013); Hele (2014); Hele and Althorpe (2016) constructing the Liouvillian, and then transforming this to path-integral normal modes. This transformation results in the observable being a function of a finite number of the lowest normal modes and truncation in the higher normal modes leads to dynamics which is both classical and conserves the QBD. Recent work has proposed a nonadiabatic Matsubara (NA-Mats) method, but to our knowledge, this does not conserve the QBD.Chowdhury and Huo (2021); Richardson et al. (2017) This is unanticipated as the Matsubara dynamics approach to the nuclear DoF is utilised, so one would expect that it may inherit the same properties. However, the electronic DoF are left unchanged in their MMST variable form. This leaves the question of if additionally following the Matsubara approach with electronic normal modes will lead to QBD conservation for a NA-Mats method.
@@
It would be highly desirable if a non-adiabatic version of Matsubara dynamics could be derived which could be proven to conserve the QBD. Although the dynamics may be too costly to evaluate for large systems (due to a probable phase factor), approximations to this may lead to accurate and inexpensive methods, in the same way that approximations to single-surface Matsubara dynamics lead to RPMD, Centroid Molecular Dynamics (CMD) and Thermostatted (T)-RPMD.Craig and Manolopoulos (2004); Ceriotti et al. (2010); Hele et al. (2015b); Hele (2016); Hele et al. (2015a); Jang and Voth (1999a, b); Hele and Suleimanov (2015)
@@
In this work, we aim to explore the properties of electronic normal modes to better understand them and to determine whether the conservation of the QBD can be obtained. We will do this by considering an electronic-only system and taking normal modes of the MMST representation, as this form is used in the the proposed NA-Mats.Chowdhury and Huo (2021) The MMST mapping Hamiltonian is known to satisfy some of our criteria, such as replicating Rabi oscillations and having accurate, classical scaling algorithms for propagation.Cook et al. (2023)
@@
The structure of this article is as follows. In section II, we present background theory, including correlation functions and Matsubara dynamics. In section III, we present the methodology of electronic normal modes of the MMST variables. Algebraically, we determine that truncating in MMST normal modes will not conserve the QBD for a single trajectory. In section IV, we present the computational results of MMST normal modes for an electronic-only system. We find that all MMST normal modes are required to obtain reasonable dynamics, and that there is no constraint on the higher normal modes. Truncating in normal modes appears to conserve the distribution for an ensemble of trajectories due to an averaging effect. We conclude in section V.
@@ II.1 Single Surface
where is the QBD, and are operators where . The related Kubo-transformed (KT) CF has a higher degree of symmetry than the standard CF in Eqn. (2),Kubo (1966); Craig and Manolopoulos (2004); Willatt (2017)
which satisfies detailed balance and reality.Craig and Manolopoulos (2004); Willatt (2017) Real time-reversal symmetry will be satisfied for both the KT-CF and standard CF when the operators are functions of only position.Willatt (2017); Craig and Manolopoulos (2004) One can obtain the standard CF from the KT-CF (and vice-versa) through a Fourier Transform, which makes it advantageous to calculate the KT-CF due to this higher degree of symmetry.Craig and Manolopoulos (2004)
When evaluating a CF in ring-polymer co-ordinates, a discretised imaginary-time path-integral version of the KT-CF is used to describe the evolution of the path-integral as a whole. Each bead is then a slice of the path-integral in imaginary time, and cyclic evaluation of this generalised KT-CF around the ring-polymer is an approximation to the whole path-integral. The generalised KT-CF is,Hele et al. (2015a); Hele and Althorpe (2013a, b); Althorpe and Hele (2013)
where the operators act once in the path-integral such that and likewise for , the indices are cyclic () and . The generalised KT-CF is then equivalent to the KT-CF in the limit when evaluating certain linear operators.Hele and Althorpe (2013a); Willatt (2017); Hele et al. (2015a) Matsubara dynamics approximates the generalised KT-CF through truncation of the Liouvillian which is discussed in further detail later.Hele et al. (2015a)
@@
and and likewise for and .Hele et al. (2015a); Hele and Ananth (2016) The form of and are different, where the latter is a simple bead-average of the Wigner transform and the former is more complex.
@@
Matsubara modes have been utilized in path-integral descriptions of equilibrium properties for many years, and gives rise to an approximate ring-polymer expression for the zero-time value of Eqn. (3).Ceperley (1995); Chakravarty (1997); Chakravarty, Gordillo, and Ceperley (1998); Freeman and Doll (1984) This means that many observables can be written as a function of only a finite number of the lowest normal modes of the nuclear position and/or momenta.Hele et al. (2015a); Hele and Althorpe (2016) The Matsubara modes have a special property where any superposition of them results in a smooth and differentiable distribution with respect to imaginary time.Hele et al. (2015a) The Liouvillian can also be expressed in terms of the normal modes, and can be truncated to be only the lowest normal modes. This is an approximation as the propagation of the higher and lower normal modes is generally coupled. Truncation of the Liouvillian in the higher normal modes surprisingly results in classical dynamics, occurring when the Matsubara modes are decoupled from the non-Matsubara modes.Hele et al. (2015a) The (unpropagated) higher normal modes can then be integrated out due to the QBD. This also conserves the QBD, as the remaining lowest path-integral modes are a smooth function of imaginary time such that Noether’s theorem can be used to prove quantum Boltzmann conservation.Hele et al. (2015a); Goldstein (1980)
II.2 Non-Adiabatic Systems
The previously proposed NA-Mats method uses this Matsubara approach for the nuclear DoF, but leaves the electronic DoF unchanged in the MMST mapping representation, and this generally does not conserve the QBD.Chowdhury and Huo (2021) We suspect that to achieve conservation of the QBD for a non-adiabatic Matsubara method, a similar approach to Matsubara dynamics for the normal modes of the electronic degrees of freedom may be required. When considering electronic normal modes, there are many metrics that one can choose from for the state populations. For example, the occupancy calculated from Meyer–Miller–Stock–Thoss (MMST) mapping variables,Meyer and Miller (1979); Stock and Thoss (2005) spin-mapping or the action-angle approach,Runeson and Richardson (2019, 2020); Meyer and Miller (1979) and it is not immediately apparent which metric to use. We ideally would like a metric that satisfies the following,Hele et al. (2015a, b)
@@ II.2 Non-Adiabatic Systems
The distribution of the higher normal modes is constrained (such that they can be integrated out when dynamics is truncated)
@@
For a non-adiabatic system, the Hamiltonian for an -level system in the diabatic representation is,
@@ III Methodology
which is the third term of Eqn. (13), but the potential matrix is independent of nuclear position as we only have an electronic system. It is important to note that the electronic position and momenta are not themselves physical observables, as they are the mapping variables for the quantum subsystem. The electronic-only Liouvillian is the electronic part of the decoupled Liouvillian in Eqn. (15),
@@
and and likewise for .Hele et al. (2015a); Hele and Ananth (2016) Again, the operators are only evaluated once in the path-integral loop such that, and likewise for .
@@ III.1 Electronic Normal Modes
where k is the normal mode index and n is the electronic state index, and we sum over j beads. The zeroth mode is related to the centroid as in Eqn. (11), the first mode is crossing from one state to another and back again once and so on. We use the same transformation matrix as for normal modes of a free ring-polymer, Eqn. (8), which is similar to the transformation obtaining Matsubara modes of the nuclear variables in Matsubara dynamics.Hele et al. (2015a); Ceriotti et al. (2010) We wish to make it clear that here we are calculating electronic normal modes and not Matsubara modes which are defined with an additional factor of .Hele et al. (2015a) We use this similar transformation as we wish to have a consistent transformation for both electronic and nuclear variables. By evaluating Hamilton’s equations of motion we can obtain the exact electronic propagation equations.Cook et al. (2023); Church et al. (2018); Richardson et al. (2017) For a single bead,
@@
which in Appendix A, we prove is also conserved in the case of decoupled nuclear and electronic DoF.
@@ III.2 The Electronic-Only Liouvillian with Normal Modes
which we show results in consistent propagation equations with Eqn. (29).
A very similar derivation can be done when there is nuclear dependence to obtain Eqn. (C.4), as seen in Appendix C. However, we note that this is not the full coupled Liouvillian but only the electronic term. We have shown that the evolution in normal modes is very similar to the evolution in beads. In the case where there is no nuclear dependence, the evolution equations are identical and evolution in normal modes is separable. However, when there is nuclear dependence, we see the slightly more complex evolution equation in Eqn.(C.4) where the potential matrix is replaced by a tensor.

### v2

# On the electronic path integral normal modes of the Meyer–Miller–Stock–Thoss representation of nonadiabatic dynamics

@@ Abstract
Accurate and efficient simulation of nonadiabatic dynamics is highly desirable for understanding charge and energy transfer in complex systems. A key criterion for obtaining an accurate method is conservation of the Quantum Boltzmann Distribution (QBD). For a single surface, Matsubara dynamics is known to conserve the QBD, as a consequence of truncating the dynamics in the higher normal modes of the imaginary-time path integral. Recently, a nonadiabatic Matsubara (NA-Mats) dynamics has been proposed (J. Chem. Phys., 2021, 154, 124124) which truncates in the normal modes of the nuclear variables but not in the electronic variables, which are described with the Meyer–Miller–Stock–Thoss (MMST) representation. Surprisingly, this NA-Mats method does not appear to conserve the QBD for a general system. This poses the question of the effect of truncating the higher path integral normal modes of the electronic variables in the MMST representation. In this article, we present what we believe is the first study of electronic normal modes of the MMST representation. We find that observables are not usually a function of a finite number of normal modes and that the higher normal modes are not constrained by the distribution, unlike in conventional nuclear normal modes. Furthermore, truncating the dynamics in MMST normal modes leads to inaccurate correlation functions and while the QBD appears conserved for an ensemble of trajectories, it is not for a single trajectory. Overall, this suggests that MMST path integral normal modes are not optimal for obtaining an accurate, QBD conserving nonadiabatic dynamics method.
@@ I Introduction
Many of the interesting and complex phenomena observed from charge and energy-transfer mechanisms occur via nonadiabatic dynamics, where the electronic and nuclear degrees of freedom are coupled. Examples include photoisomerism, ultrafast relaxation via intersystem crossing in DNA or charge carriers in solar cell materials, and electron transfers in various photosynthetic systems and catalytic materials.Polli et al. (2010); Wan et al. (1999); Hammes-Schiffer and Stuchebrukhov (2010); Marcus and Sutin (1985); Zhu and Podzorov (2015); Cheng and Fleming (2009); Domcke and Yarkony (2012) As these processes often occur on very short timescales,Mukherjee et al. (2025) it can be experimentally challenging to understand the dynamics.
Hence, many theoretical methods have been developed to simulate nonadiabatic dynamics.Althorpe et al. (2016); Nelson et al. (2020) These range from full quantum methods such as Multi-Configurational Time-Dependent Hartree-Fock (MCTDH) and wavepacket propagation,Beck et al. (2000); Van Haeften, Ash, and Worth (2023); Wang and Thoss (2003); Shin and Metiu (1996) to more approximate semiclassical methods including initial-value representations (IVR),Stock and Thoss (1997); Miller (1970); Wang, Sun, and Miller (1998); Sun, Wang, and Miller (1998a, b); Church et al. (2018); Ananth, Venkataraman, and Miller (2007) nonadiabatic ring-polymer molecular dynamics (RPMD) and their derivatives.Richardson and Thoss (2013); Ananth (2013); Hele (2011); Hele et al. (2015a); Cao and Voth (1993, 1994a, 1994b, 1994c); Rossi, Ceriotti, and Manolopoulos (2014) Another useful class of methods are mixed quantum-classical methods, including surface hopping, Ehrenfest dynamics, and mapping-based methods.Kapral (2016); Tully and Preston (1971); Tully (2012); Shakib and Huo (2017); Shalashilin (2011); Zimmermann and Vaníček (2014); Meyer and Miller (1979); Stock and Thoss (1997, 2005); Runeson and Richardson (2019); Mannouch and Richardson (2023) These methods often have a good balance between computational cost and accuracy.
@@
When considering the development of a nonadiabatic dynamics method, it would be highly desirable if the method satisfied the following criteria:
@@
While many of the existing methods satisfy many of these criteria, as far as we are aware, none satisfy all of them simultaneously and exactly for a general quantum system.
@@
Conservation of the QBD is a particularly interesting criterion to investigate, as for single-surface dynamics, Matsubara dynamics conserves the QBD.Willatt (2017); Hele et al. (2015a) Matsubara dynamics achieves this by starting with an (exact quantum) Generalized Kubo Transform,Hele and Althorpe (2013a, b); Althorpe and Hele (2013); Hele (2014); Hele and Althorpe (2016) constructing the Liouvillian, and then transforming this to path integral normal modes. This transformation results in the observable being a function of a finite number of the lowest normal modes and truncation in the higher normal modes leads to dynamics which is both classical and conserves the QBD. Recent work has proposed a nonadiabatic Matsubara (NA-Mats) method, but to our knowledge, this does not conserve the QBD.Chowdhury and Huo (2021); Richardson et al. (2017) This is unanticipated as the Matsubara dynamics approach to the nuclear DoF is utilised, so one would expect that it may inherit the same properties. However, the electronic DoF are left unchanged in their MMST variable form. This leaves the question of if additionally following the Matsubara approach with electronic normal modes will lead to QBD conservation for a NA-Mats method.
@@
It would be highly desirable if a nonadiabatic version of Matsubara dynamics could be derived which could be proven to conserve the QBD. Although the dynamics may be too costly to evaluate for large systems (due to a probable phase factor), approximations to this may lead to accurate and inexpensive methods, in the same way that approximations to single-surface Matsubara dynamics lead to RPMD, Centroid Molecular Dynamics (CMD) and Thermostatted (T)-RPMD.Craig and Manolopoulos (2004); Ceriotti et al. (2010); Hele et al. (2015b); Hele (2016); Hele et al. (2015a); Jang and Voth (1999a, b); Hele and Suleimanov (2015)
@@
In this work, we aim to explore the properties of electronic normal modes to better understand them and to determine whether the conservation of the QBD can be obtained. We will do this by considering an electronic-only system and take normal modes of the MMST representation, as this form is used in the the proposed NA-Mats.Chowdhury and Huo (2021) The MMST mapping Hamiltonian is known to satisfy some of our criteria, such as replicating Rabi oscillations and having accurate, classical scaling algorithms for propagation.Cook et al. (2023)
@@
The structure of this article is as follows. In section II, we present background theory, including correlation functions and Matsubara dynamics. In section III, we present the methodology of electronic normal modes of the MMST variables. Algebraically, we determine that truncating in MMST normal modes will not conserve the QBD for a single trajectory. In section IV, we present the computational results of MMST normal modes for an electronic-only system. We find that all MMST normal modes are required to obtain reasonable dynamics, and that there is no narrowing in the distribution of the higher normal modes. Truncating in normal modes appears to conserve the distribution for an ensemble of trajectories due to an averaging effect. We conclude in section V.
@@ II.1 Single Surface
where is the QBD, and are operators where and is the partition function. The related Kubo-transformed (KT) CF has a higher degree of symmetry than the standard CF in Eqn. (2),Kubo (1966); Craig and Manolopoulos (2004); Willatt (2017)
which satisfies detailed balance and reality.Craig and Manolopoulos (2004); Willatt (2017) In addition, (time-reversal symmetry) will be satisfied for both the KT-CF and standard CF when the operators are functions of only position.Willatt (2017); Craig and Manolopoulos (2004) One can obtain the standard CF from the KT-CF (and vice-versa) through a Fourier Transform, which makes it advantageous to calculate the KT-CF due to this higher degree of symmetry.Craig and Manolopoulos (2004)
When evaluating a CF in ring-polymer co-ordinates, a discretised imaginary-time path integral version of the KT-CF is used to describe the evolution of the path integral as a whole. Each bead is then a slice of the path integral in imaginary time, and cyclic evaluation of this generalised KT-CF around the ring-polymer is an approximation to the whole path integral. The generalised KT-CF is,Hele et al. (2015a); Hele and Althorpe (2013a, b); Althorpe and Hele (2013)
where the operators act once in the path integral such that and likewise for , the indices are cyclic () and . The generalised KT-CF is then equivalent to the KT-CF in the limit when the operators can be written as a linear sum of operators that act once in the path integral, as is defined above.Hele and Althorpe (2013a); Willatt (2017); Hele et al. (2015a) For example, the flux and side operators utilised in rate theory are non-linear under this definition.Althorpe and Hele (2013); Hele and Althorpe (2013a, b); Althorpe and Hele (2013); Hele (2014) Matsubara dynamics approximates the generalised KT-CF through truncation of the Liouvillian which is discussed in further detail later.Hele et al. (2015a)
@@
where is the symmetric operator, is the time-evolved operator and, and likewise for and .Hele et al. (2015a); Hele and Ananth (2016) The form of and are different, where the latter is a simple bead-average of the Wigner transform and the former is more complex.
@@
Matsubara modes have been utilized in path integral descriptions of equilibrium properties for many years, and gives rise to an approximate ring-polymer expression for the zero-time value of Eqn. (3).Ceperley (1995); Chakravarty (1997); Chakravarty, Gordillo, and Ceperley (1998); Freeman and Doll (1984) This means that many observables can be written as a function of only a finite number of the lowest normal modes of the nuclear position and/or momenta.Hele et al. (2015a); Hele and Althorpe (2016) The Matsubara modes have a special property where any superposition of them results in a smooth and differentiable distribution with respect to imaginary time.Hele et al. (2015a) The Liouvillian can also be expressed in terms of the normal modes, and can be truncated to be only the lowest normal modes. This is an approximation as the propagation of the higher and lower normal modes is generally coupled. Truncation of the Liouvillian in the higher normal modes surprisingly results in classical dynamics, occurring when the Matsubara modes are decoupled from the non-Matsubara modes.Hele et al. (2015a) The (unpropagated) higher normal modes can then be integrated out due to the QBD. This also conserves the QBD, as the remaining lowest path integral modes are a smooth function of imaginary time such that Noether’s theorem can be used to prove quantum Boltzmann conservation.Hele et al. (2015a); Goldstein (1980)
II.2 Nonadiabatic Systems
The previously proposed NA-Mats method uses this Matsubara approach for the nuclear DoF, but leaves the electronic DoF unchanged in the MMST mapping representation, and this generally does not conserve the QBD.Chowdhury and Huo (2021) We suspect that to achieve conservation of the QBD for a nonadiabatic Matsubara method, a similar approach to Matsubara dynamics for the normal modes of the electronic degrees of freedom may be required. When considering electronic normal modes, there are many metrics that one can choose from for the state populations. For example, the occupancy calculated from Meyer–Miller–Stock–Thoss (MMST) mapping variables,Meyer and Miller (1979); Stock and Thoss (2005) spin-mapping or the action-angle approach,Runeson and Richardson (2019, 2020); Meyer and Miller (1979) and it is not immediately apparent which metric to use. We ideally would like a metric that satisfies the following,Hele et al. (2015a, b)
@@ II.2 Nonadiabatic Systems
The distribution of the higher normal modes narrows (such that they can be integrated out when dynamics is truncated)
@@
For a nonadiabatic system, the Hamiltonian for an -level system in the diabatic representation is,
@@ III Methodology
which is the third term of Eqn. (13), but the potential matrix is independent of nuclear position as we only have an electronic system. It is important to note that the electronic position and momenta are not themselves physical observables, as they are the mapping variables for the quantum subsystem. The electronic-only Liouvillian is the electronic part of the decoupled Liouvillian in Eqn. (II.2),
@@
where is the symmetric operator, is the time-evolved operator and, and likewise for .Hele et al. (2015a); Hele and Ananth (2016) Again, the operators are linear in the sense that and likewise for , such that they only act once in the path integral loop. However, as the operators are not a function of only position, may not necessarily be equal to for a general (complex) nonadiabatic Hamiltonian.
@@ III.1 Electronic Normal Modes
where k is the normal mode index and n is the electronic state index, and we sum over j beads. The zeroth mode is related to the centroid as in Eqn. (11). We use the same transformation matrix as for normal modes of a free ring-polymer, Eqn. (8), which is similar to the transformation obtaining Matsubara modes of the nuclear variables in Matsubara dynamics.Hele et al. (2015a); Ceriotti et al. (2010) We wish to make it clear that here we calculate electronic normal modes and not Matsubara modes which are defined with an additional factor of .Hele et al. (2015a) We use this similar transformation as we wish to have a consistent transformation for both electronic and nuclear variables. We note that as the electronic-only system is described by a set of uncoupled harmonic oscillators at the same frequency, transforming into normal modes does not change the dynamics.
By evaluating Hamilton’s equations of motion we can obtain the exact electronic propagation equations.Cook et al. (2023); Church et al. (2018); Richardson et al. (2017) For a single bead,
@@
which we show in Appendix A is also conserved in the case of decoupled nuclear and electronic DoF.
@@ III.2 The Electronic-Only Liouvillian with Normal Modes
which results in consistent propagation equations with Eqn. (29).
A very similar derivation can be done when there is nuclear dependence to obtain Eqn. (59), as seen in Appendix C. However, we note that this is not the full coupled Liouvillian but only the electronic term. We have shown that the evolution in normal modes is very similar to the evolution in beads. In the case where there is no nuclear dependence, the evolution equations are identical and evolution in normal modes is separable. However, when there is nuclear dependence, we see the slightly more complex evolution equation in Eqn.(59) where the potential matrix is replaced by a tensor.